OpenStudy (anonymous):

A boat requires four hours to travel a given distance upstream against a five mile per hour current, but only requires three hours to travel downstream with that same current back to the dock. Find the speed of the boat in still water.

OpenStudy (anonymous):

Upstream against the current = v - 5 Downstream with the current = v + 5 Time traveling upstream = 4 hrs Time traveling downstream = 3 hrs 4(v - 5) = 3(v + 5) 4v - 20 = 3v + 15 - 3v - 3v v - 20 = 15 + 20 +20 v = 35 The speed of the boat in still water is 35 mph

OpenStudy (anonymous):

Do a google search, it'll help you alot.